# Towards Synthetic Magnetic Turbulence with Coherent Structures

### v1

@@ Abstract
Synthetic turbulence is a relevant tool to study complex astrophysical and space plasma environments inaccessible by direct simulation. However, conventional models lack intermittent coherent structures, which are essential in realistic turbulence. We present a novel method, featuring coherent structures, conditional structure function scaling and fieldline curvature statistics comparable to magnetohydrodynamic turbulence. Enhanced transport of charged particles is investigated as well. This method presents significant progress towards physically faithful synthetic turbulence.
synthetic turbulence — MHD turbulence — cosmic rays — solar energetic particles — intermittency — coherent structures
@@
In this letter we present progress towards a model for synthetic magnetic turbulence featuring intermittent coherent structures. We implement the model as a fast algorithm, which produces a random three-dimensional divergence-free vector field, resembling a turbulent magnetic field $\mathbf{b}(\mathbf{x})$  [^1] The code is publicly available at https://github.com/jerluebke/synth-mag-turb and archived at doi:10.5281/zenodo.10515965. The model is a combination of a continuous cascade [27] and the minimal multiscale Lagrangian map [22, 23]. Additionally, we propose a set of quantities to assess the physical fidelity of synthetic turbulence models, consisting of the energy spectrum, conditional structure function scaling, the fieldline curvature distribution and running diffusion coefficients of charged test particles. Based on these quantities, we compare the proposed synthetic turbulence model with an incompressible resistive MHD turbulence simulation and an intermittent synthetic turbulence model without coherent structures. We also consider the phase-randomized counterparts of the three turbulence models to account for differences in the energy spectra. We conclude by explaining the shape of the MHD fieldline curvature distribution by means of a weighted sum of Gaussian components.
@@
Methods — We start by extending the continuous cascade in wavelet space [27] to three-dimensional divergence-free vector fields. The continuous cascade at scale $l$ and position $\mathbf{x}$ is represented by a log-infinitely divisible process $e^{\omega_{l}(\mathbf{x})}$, which gives the scale- and position-dependent intensity of a vector field $\mathbf{v}(\mathbf{x})$. This field is obtained by a vector-valued wavelet transform of $l^{H}e^{\omega_{l}}$ over the inertial range scales $l_{\mathrm{min}}<l<l_{0}$ as
@@
The infinitely divisible process $\omega_{l}(\mathbf{x})$ is defined on cone-like subsets of the position-scale half-space $\mathbb{R}^{d}\times\mathbb{R}_{>0}$ equipped with the measure $l^{-d-1}\mathop{}\!\mathrm{d}{\mathbf{x}}\mathop{}\!\mathrm{d}{l}$ and has a cumulant generating function $\phi(q)$. Thus, the moments of the intensity process can be computed as $\langle e^{q\omega_{l}}\rangle=(l_{0}/l)^{c_{d}\phi(q)}\ \forall q\in\mathbb{N}$, where $c_{d}=2^{-d}\pi^{d/2}/\Gamma(d/2+1)$ comes from the scale-space cone volume. We consider a Gaussian distribution with $\phi(q)=\mu/2(p^{2}-p)$ with intermittency parameter $\mu$, in which case $e^{\omega_{l}}$ corresponds to the Gaussian multiplicative chaos employed by Pereira et al. [24] and related works.
As shown below, the vector field given by Equation (1) exhibits (isotropic) anomalous scaling, but lacks coherent features, so we introduce advective structures by adapting the minimal multiscale Lagrangian map (MMLM) [22, 23] to our framework. In short, the MMLM procedure considers only the advective part of the magnetic field evolution equation, i.e. $\left(\partial_{t}+\mathbf{u}\cdot\nabla\right)\mathbf{b}=0$, which can be solved at time $\tau$ with the Lagrangian ansatz $\mathbf{b}(\mathbf{x}_{\tau},\tau)=\mathbf{b}(\mathbf{x}_{0},0)$ and the linearized solution $\mathbf{x}_{\tau}=\mathbf{x}_{0}+\tau\mathbf{u}(\mathbf{x}_{0},0)$. Usually, two initially Gaussian random vector fields representing $\mathbf{u}(\mathbf{x})$ and $\mathbf{b}(\mathbf{x})$ are deformed on successively finer scales $l_{i}$ by applying the linearized Lagrangian solution with $\tau\propto l_{i}$ to the underlying regular grid, as described in the references.
In our framework, we first generate a random velocity field $\mathbf{u}(\mathbf{x})$ according to Equation (1) and accumulate the deformations of the grid $\mathbf{x}$ over all scales $l_{i}$ given by the discretization of the scale integral. Since our algorithm is formulated in terms of a vector potential, the Lagrangian step is computed with intermediate results $\mathbf{u}_{l_{i}}=\nabla\times\mathbf{a}_{l_{i}}^{\mathbf{u}}$. By using finite difference or Fourier methods to compute the curl, the velocity field $\mathbf{u}(\mathbf{x})$ remains unaffected by the deformed grid. We then generate an independent random magnetic vector potential $\mathbf{a}_{l_{\mathrm{min}}}^{\mathbf{b}}(\mathbf{x})$ with Equation (1) and evaluate it via inverse distance weighting at the deformed grid after the final step of the scale integral and before applying the curl. Thus, we are solving the advection equation for the magnetic vector potential
@@
which is exact in two dimensions and making it especially suitable for strong guide field situations [41].
@@
The deformation timescale $\tau=c\,l_{i}/\operatorname{max}(\|\mathbf{u}_{l_{i}}\|)$ is normalized to the maximal value of the current velocity magnitude and governed by the constant $c$. This constant is a free parameter, which must be chosen carefully, as it should be large enough for coherent structures to emerge, but not too large to avoid decorrelation. For higher values of $c$, energy accumulates on smaller scales, which is corrected by reweighting $\mathbf{a}^{\mathbf{b}}_{l_{\mathrm{min}}}(\mathbf{x})$ after interpolation with $k^{\delta/2}$, where $\delta$ is the deviation of the energy spectrum scaling from the expected $-5/3$rd scaling. Additionally, we mimic the effect of dissipation by applying a low-pass filter $\exp(-{k^{2}}/{2k_{0}^{2}})$.
@@
For comparison, we perform a direct numerical simulation of incompressible resistive MHD turbulence in a three-dimensional periodic box with resolution $N^{3}=1024^{3}$, no background field and equal diffusivity and resistivity $\nu=\eta=1.2\times 10^{-3}$ [42]. The velocity is driven on Fourier modes $1\leq k\leq 3$ with the random forcing proposed by Alvelius [43], which exhibits very low mean cross-helicity and low noise in the time evolution of turbulence bulk quantities. For this purpose we employed the pseudo-spectral code SpecDyn, which was developed in the context of magnetic dynamo action and is tailored for use on modern HPC systems [44, 45].
@@
Results — We generate sample fields of the continuous cascade process (CC) given by Equation (1) and its minimal multiscale Lagrangian mapping extension (LM) corresponding to Equation (2) on a three-dimensional periodic grid with resolution $N^{3}=1024^{3}$. The parameters of both processes are $H=1/3$$l_{0}=0.5$ and $l_{\mathrm{min}}\approx 1.5\mathop{}\!\mathrm{d}{x}$. The intermittency parameter of the CC process is $\mu=0.4$, and the additional parameters of the LM process are $\mu=0.1$$c=0.2$$\delta=-0.45$, and $k_{0}=256$.
@@
For visual inspection, slices of field strength and fieldline curvature are plotted in Figure 1, together with a plot of the radially averaged energy spectra. MHD turbulence is characterized by elongated and intricately intertwined coherent structures, while the CC field consists of incoherent intermittent “clouds” of large field strength values. The LM field exhibits thin and intense coherent structures, which are more spatially isolated. While CC and LM match the expected $-5/3$rd scaling rather well, the impact of resistivity of the MHD simulation becomes apparent.
@@
We employ conditional structure functions [46] to study intermittency with respect to the local structure of the field. Given a displacement vector $\mathbf{d}$ at a point $\mathbf{X}$, we consider increments $\delta\mathbf{b}_{\perp}=\mathbf{b}_{\perp}(\mathbf{X}+\mathbf{d})-\mathbf{b}_{\perp}(\mathbf{X})$ perpendicular to the local mean field $\mathbf{b}_{\mathrm{local}}=\big{(}\mathbf{b}(\mathbf{X}+\mathbf{d})+\mathbf{b}(\mathbf{X})\big{)}/2$. Averages over these increments are taken conditionally on the direction of $\mathbf{d}$ in an instantaneous local basis given by $\hat{\mathbf{b}}_{\mathrm{local}}$, the normal direction of the fluctuations $\delta\hat{\mathbf{b}}_{\perp,N}$ with $\delta\mathbf{b}_{\perp,N}=\delta\mathbf{b}_{\perp}-\big{(}\delta\mathbf{b}_{\perp}\cdot\hat{\mathbf{b}}_{\mathrm{local}}\big{)}\hat{\mathbf{b}}_{\mathrm{local}}$ and the binormal direction $\hat{\mathbf{b}}_{\mathrm{local}}\times\delta\hat{\mathbf{b}}_{\perp,N}$. Based on this, the scaling exponents of the conditional averages $\big{\langle}\|\delta\mathbf{b}_{\perp}\|^{p}|\,\hat{\mathbf{d}}\,\big{\rangle}\propto d^{\zeta_{p}}$ are denoted as $\zeta_{p,\parallel}$$\zeta_{p,\delta\hat{\mathbf{b}}}$ or $\zeta_{p,\perp}$ depending on $\mathbf{d}$ being aligned with $\hat{\mathbf{b}}_{\mathrm{local}}$$\delta\hat{\mathbf{b}}_{\perp,N}$ or $\hat{\mathbf{b}}_{\mathrm{local}}\times\delta\hat{\mathbf{b}}_{\perp,N}$.
Figure 2 shows the normalized scaling exponents $\zeta_{p}/\zeta_{3}$ as well as the raw values of $\zeta_{3}$ for the three models and the three directions. The further $\zeta_{p}/\zeta_{3}$ deviates from the linear case, the more intermittent is the process, and smaller values of $\zeta_{3}$ correspond to a rougher process. In the MHD case, the direction parallel to the local mean field is the most smooth and non-intermittent, while the fluctuation and perpendicular direction are equally intermittent, the latter being the most rough. In contrast to this, the CC case shows no significant difference between the three directions, which are all equally rough and intermittent and which is expected from the lack of coherent structures. Lastly, the LM case exhibits again directional dependency, with parallel being the most non-intermittent and perpendicular the most intermittent. However, the fluctuation direction resembles more closely the parallel than the perpendicular direction.
@@
Figure 3 shows the distributions $p(\kappa)$ for the three cases and their random-phase counterparts. The MHD case, in agreement with the literature [38, 39], behaves asymptotically as $p(\kappa)\underset{\kappa\to\infty}{\sim}\kappa^{-2.5}$. The CC case agrees, apart from being slightly wider, with the Gaussian random-phase fields, which scale distinctly as $p(\kappa)\underset{\kappa\to\infty}{\sim}\kappa^{-4}$. Finally, the LM case appears as an intermediate case between the previous two cases; its distribution function $p(\kappa)$ is slightly more narrow than the MHD case and around the slightly right-shifted peak, $p(\kappa)$ faintly resembles the $\kappa^{-2.5}$ scaling, before adjusting to the Gaussian $\kappa^{-4}$ scaling.
@@
Since the random-phase fields are Gaussian random variables, their normalized fieldline curvature follows a universal distribution $p(\kappa/\kappa_{\mathrm{peak}})$, which is independent of the energy spectrum. However, $\kappa_{\mathrm{peak}}$ does depend on the energy spectrum, e.g. via the slope $s$ in case of a power-law spectrum $\sim k^{-s}$. Flatter spectra implicate more energy on small scales, resulting in more contributions from high curvatures and consequently larger $\kappa_{\mathrm{peak}}$. This connection is illustrated in Figures 4, where $p_{\mathrm{MHD}}(\kappa)$ is modeled as a weighted sum of Gaussian components, similar to the description of the turbulent velocity increment distribution function as a Gaussian scale mixture [47].
@@
Conclusion — We have presented a novel algorithm for synthetic magnetic turbulence based on a combination of the continuous cascade model, generalized to three-dimensional divergence-free vector fields, and the minimal multiscale Lagrangian map. We compare this algorithm with an incompressible resistive MHD turbulence simulation and the pure three-dimensional continuous cascade model, which is intermittent but lacks coherent structures. This comparison is done by means of visual inspection, the energy spectrum, conditional structure function scaling, the fieldline curvature distribution and running diffusion coefficients of charged test particles.

### v3

# Institut für Theoretische Physik IV — Ruhr-Universität Bochum, 44801 Bochum, Germany

@@ Abstract
Synthetic turbulence is a relevant tool to study complex astrophysical and space plasma environments inaccessible by direct simulation. However, conventional models lack intermittent coherent structures, which are essential in realistic turbulence. We present a novel method featuring coherent structures, conditional structure function scaling and fieldline curvature statistics comparable to magnetohydrodynamic turbulence. Enhanced transport of charged particles is investigated as well. This method presents significant progress towards physically faithful synthetic turbulence.
1 Introduction
@@
In this letter we present progress towards a model for synthetic magnetic turbulence featuring intermittent coherent structures. We implement the model as a fast algorithm, which produces a random three-dimensional divergence-free vector field, resembling a turbulent magnetic field $\mathbf{b}(\mathbf{x})$  [^1] The code is publicly available at https://github.com/jerluebke/synth-mag-turb and archived at doi:10.5281/zenodo.10515965. A pseudocode listing is provided as supplementary material.. The model is a combination of a continuous cascade [27] and the minimal multiscale Lagrangian map [22, 23]. Additionally, we propose a set of quantities to assess the physical fidelity of synthetic turbulence models, consisting of the energy spectrum, conditional structure function scaling, the fieldline curvature distribution and running diffusion coefficients of charged test particles. Based on these quantities, we compare the proposed synthetic turbulence model with an incompressible resistive MHD turbulence simulation and an intermittent synthetic turbulence model without coherent structures. We also consider the phase-randomized counterparts of the three turbulence models to account for differences in the energy spectra. We conclude by explaining the shape of the MHD fieldline curvature distribution by means of a weighted sum of Gaussian components.
@@
2 Methods
We start by extending the continuous cascade in wavelet space [27] to three-dimensional divergence-free vector fields. The continuous cascade at scale $l$ and position $\mathbf{x}$ is represented by a log-infinitely divisible process $e^{\omega_{l}(\mathbf{x})}$, which gives the scale- and position-dependent intensity of a vector field $\mathbf{v}(\mathbf{x})$. This field is obtained by a vector-valued wavelet transform of $l^{H}e^{\omega_{l}}$ over the inertial range scales $l_{\mathrm{min}}<l<l_{0}$ as
@@ 2 Methods
The infinitely divisible process $\omega_{l}(\mathbf{x})$ is defined on cone-like subsets of the position-scale half-space $\mathbb{R}^{d}\times\mathbb{R}_{>0}$ equipped with the measure $l^{-d-1}\mathop{}\!\mathrm{d}{\mathbf{x}}\mathop{}\!\mathrm{d}{l}$ and has a cumulant generating function $\phi(q)$. Thus, the moments of the intensity process can be computed as $\langle e^{q\omega_{l}}\rangle=(l_{0}/l)^{c_{d}\phi(q)}\ \forall q\in\mathbb{N}$, where $c_{d}=2^{-d}\pi^{d/2}/\Gamma(d/2+1)$ comes from the scale-space cone volume. We consider a Gaussian distribution with $\phi(q)=\mu/2(p^{2}-p)$ with intermittency parameter $\mu$, in which case $e^{\omega_{l}}$ corresponds to the Gaussian multiplicative chaos employed in [24] and related works.
As shown below, the vector field given by Equation (1) exhibits (isotropic) anomalous scaling, but lacks coherent features, so we introduce advective structures by adapting the minimal multiscale Lagrangian map (MMLM) for MHD [23] to our framework. In short, the MMLM procedure considers only the advective part of the magnetic field evolution equation, i.e. $\left(\partial_{t}+\mathbf{u}\cdot\nabla\right)\mathbf{b}=0$, which can be solved at time $\tau$ with the Lagrangian ansatz $\mathbf{b}(\mathbf{x}_{\tau},\tau)=\mathbf{b}(\mathbf{x}_{0},0)$ and the linearized solution $\mathbf{x}_{\tau}=\mathbf{x}_{0}+\tau\mathbf{u}(\mathbf{x}_{0},0)$. Usually, two initially Gaussian random vector fields representing $\mathbf{u}(\mathbf{x})$ and $\mathbf{b}(\mathbf{x})$ are deformed on successively finer scales $l_{i}$ by applying the linearized Lagrangian solution with $\tau\propto l_{i}$ to the underlying regular grid, as described in the references.
In our framework, we generate two independent random fields according to Equation (1), which are represented as vector potentials and represent the velocity field $\mathbf{u}(\mathbf{x})=\nabla\times\mathbf{a}^{\mathbf{u}}(\mathbf{x})$ and the magnetic field $\mathbf{b}(\mathbf{x})=\nabla\times\mathbf{a}^{\mathbf{b}}(\mathbf{x})$. When generating $\mathbf{u}(\mathbf{x})$ scale-by-scale, we make use of the discretization of the scale integral in Equation (1) as a sum $\int_{l_{\mathrm{min}}}^{l_{0}}\cdots\mathop{}\!\mathrm{d}{l}\approx\sum_{l_{i}=l_{0}}^{l_{\mathrm{min}}}\cdots\Delta{l}$, which goes from large to small scales. Specifically, we accumulate the deformations of the grid $\mathbf{x}$ by the intermediate results $\mathbf{u}_{l_{i}}=\nabla\times\mathbf{a}^{\mathbf{u}}_{l_{i}}$ over all scales $l_{i}$. We then interpolate the random magnetic vector potential $\mathbf{a}^{\mathbf{b}}(\mathbf{x})$ at the final deformed grid. Note that this interpolation is not done intertwined with the scale-by-scale generation of $\mathbf{a}^{\mathbf{b}}(\mathbf{x})$, but only once at the end of the procedure. Thus, we are solving the advection equation for the magnetic vector potential
@@
which is exact in two dimensions making this approach especially suitable for strong guide field situations [40]. Herein lies the key difference with the previous MMLM procedure, which applied the interpolation directly to the field $\mathbf{b}(\mathbf{x})$ in a scale-by-scale fashion. It should further be noted, that the curl of the intermediate results $\mathbf{u}_{l_{i}}=\nabla\times\mathbf{a}^{\mathbf{u}}_{l_{i}}$ is computed on a uniform grid, so while $\mathbf{u}_{l_{i}}$ serves as an indicator for the grid deformation, it is not affected by it.
@@
The deformation timescale $\tau=c\,l_{i}/\operatorname{max}(\|\mathbf{u}_{l_{i}}\|)$ is normalized to the maximal value of the current velocity magnitude and governed by the constant $c$. This constant is a free parameter, which must be chosen carefully, as it should be large enough for coherent structures to emerge, but not too large to avoid decorrelation. For higher values of $c$, energy accumulates on smaller scales, which is corrected by reweighting $\mathbf{a}^{\mathbf{b}}_{l_{\mathrm{min}}}(\mathbf{x})$ after interpolation with $k^{\delta/2}$, where $\delta$ is the deviation of the energy spectrum scaling from the expected $-5/3$rd scaling. Additionally, we mimic the effect of dissipation by applying a low-pass filter $\exp(-{k^{2}}/{2k_{0}^{2}})$ controlled by the artificial dissipation wavenumber $k_{0}$.
@@
For comparison, we perform a direct numerical simulation of incompressible resistive MHD turbulence in a three-dimensional periodic box with resolution $N^{3}=1024^{3}$, no background field and equal diffusivity and resistivity $\nu=\eta=1.2\times 10^{-3}$ [41]. We obtain Taylor-scale Reynolds numbers $\mathcal{R}_{\lambda}=439$ for the velocity field and $\mathcal{R}_{\lambda,m}=94$ for the magnetic field. The velocity is driven on Fourier modes $1\leq k\leq 3$ with the random forcing proposed in [42], which exhibits very low mean cross-helicity and low noise in the time evolution of turbulence bulk quantities. For this purpose we employed the pseudo-spectral code SpecDyn, which was developed in the context of magnetic dynamo action and is tailored for use on modern HPC systems [43, 44].
3 Results
@@
We generate sample fields of the continuous cascade process (CC) given by Equation (1) and its minimal multiscale Lagrangian mapping extension (LM) corresponding to Equation (2) on a three-dimensional periodic grid with resolution $N^{3}=1024^{3}$. The parameters of both processes are $H=1/3$$l_{0}=0.5$ and $l_{\mathrm{min}}\approx 1.5\mathop{}\!\mathrm{d}{x}$. The intermittency parameter of the CC process is $\mu=0.4$, and the additional parameters of the LM process are $\mu=0.1$$c=0.2$$\delta=-0.45$, and $k_{0}=256$.
@@
For visual inspection, slices of field strength and fieldline curvature are plotted in Figure 1, together with a plot of the radially averaged energy spectra. MHD turbulence is characterized by elongated and intricately intertwined coherent structures, while the CC field consists of incoherent intermittent “clouds” of large field strength values. The LM field exhibits thin and intense coherent structures, which are more spatially isolated. The CC spectrum matches the expected $-5/3$rd scaling well, the LM spectrum exhibits a bottleneck effect at high wavenumbers, and the MHD spectrum is affected by strong resistivity. The bottleneck effect of the LM spectrum is due to the simple advective nature of the Lagrangian mapping, which causes accumulation of energy on small scales. This is counteracted by reweighting the vector potential with $k^{\delta/2}$ and applying artificial dissipation, however, when balancing the relevant parameters $c$$\delta$ and $k_{0}$, a residual bottleneck effect remains. In order to take potential influences of these different spectra into account, the analysis of curvature and particle transport is also performed with the respective phase-randomized fields.
@@
We employ conditional structure functions [45] to study intermittency with respect to the local structure of the field. See also [46] for a discussion of local anisotropy statistics. Given a displacement vector $\mathbf{d}$ at a point $\mathbf{X}$, we consider increments $\delta\mathbf{b}_{\perp}=\mathbf{b}_{\perp}(\mathbf{X}+\mathbf{d})-\mathbf{b}_{\perp}(\mathbf{X})$ perpendicular to the local mean field $\mathbf{b}_{\mathrm{local}}=\big{(}\mathbf{b}(\mathbf{X}+\mathbf{d})+\mathbf{b}(\mathbf{X})\big{)}/2$. Averages over these increments are taken conditionally on the direction of $\mathbf{d}$ in an instantaneous local basis given by $\hat{\mathbf{b}}_{\mathrm{local}}$, the normal direction of the fluctuations $\delta\hat{\mathbf{b}}_{\perp,N}$ with $\delta\mathbf{b}_{\perp,N}=\delta\mathbf{b}_{\perp}-\big{(}\delta\mathbf{b}_{\perp}\cdot\hat{\mathbf{b}}_{\mathrm{local}}\big{)}\hat{\mathbf{b}}_{\mathrm{local}}$ and the binormal direction $\hat{\mathbf{b}}_{\mathrm{local}}\times\delta\hat{\mathbf{b}}_{\perp,N}$. Based on this, the scaling exponents of the conditional averages $\big{\langle}\|\delta\mathbf{b}_{\perp}\|^{p}|\,\hat{\mathbf{d}}\,\big{\rangle}\propto d^{\zeta_{p}}$ are denoted as $\zeta_{p,\parallel}$$\zeta_{p,\delta\hat{\mathbf{b}}}$ or $\zeta_{p,\perp}$ depending on $\mathbf{d}$ being aligned with $\hat{\mathbf{b}}_{\mathrm{local}}$$\delta\hat{\mathbf{b}}_{\perp,N}$ or $\hat{\mathbf{b}}_{\mathrm{local}}\times\delta\hat{\mathbf{b}}_{\perp,N}$.
Figure 2 shows the normalized scaling exponents $\zeta_{p}/\zeta_{3}$ as well as the raw values of $\zeta_{3}$ for the three models and the three directions. The further $\zeta_{p}/\zeta_{3}$ deviates from the linear case, the more intermittent is the process, and smaller values of $\zeta_{3}$ correspond to a rougher process. In the MHD case, the field is the most smooth and non-intermittent parallel to the local mean field direction $\hat{\mathbf{\mathbf{b}}}_{\mathrm{local}}$, while it is equally intermittent in the direction of fluctuations $\delta\hat{\mathbf{\mathbf{b}}}_{\perp,N}$ and the binormal direction $\hat{\mathbf{\mathbf{b}}}_{\mathrm{local}}\times\delta\hat{\mathbf{\mathbf{b}}}_{\perp,N}$, and being the most rough in the binormal direction. In contrast to this, the CC field shows no significant difference between the three directions, which are all equally rough and intermittent, which is expected from the lack of coherent structures. Lastly, the LM field exhibits again directional dependency, with the parallel direction $\hat{\mathbf{\mathbf{b}}}_{\mathrm{local}}$ being the most non-intermittent and the binormal direction $\hat{\mathbf{\mathbf{b}}}_{\mathrm{local}}\times\delta\hat{\mathbf{\mathbf{b}}}_{\perp,N}$ being the most intermittent. However, the field is in the fluctuation direction $\delta\hat{\mathbf{\mathbf{b}}}_{\perp,N}$ less intermittent compared to the MHD case.
@@
Figure 3 shows the distributions $p(\kappa)$ for the three cases and their random-phase counterparts. The MHD case, in agreement with the literature [38, 39], behaves asymptotically as $p(\kappa)\underset{\kappa\to\infty}{\sim}\kappa^{-2.5}$. Note, as shown by [38], the scaling becomes $\kappa^{-2}$ in 2D turbulence, while similar values are expected for general collisionless plasmas. The CC case agrees, apart from being slightly wider, with the Gaussian random-phase fields, which scale distinctly as $p(\kappa)\underset{\kappa\to\infty}{\sim}\kappa^{-4}$. Finally, the LM case appears as an intermediate case between the previous two cases; its distribution function $p(\kappa)$ is slightly more narrow than the MHD case and around the slightly right-shifted peak, $p(\kappa)$ faintly resembles the $\kappa^{-2.5}$ scaling, before adjusting to the Gaussian $\kappa^{-4}$ scaling.
@@
Since the random-phase fields are Gaussian random variables, we expect a universal normalized fieldline curvature distribution $p(\kappa/\kappa_{\mathrm{peak}})$, which is independent of the energy spectrum. This behaviour is observed in numerical experiments and also indicated by [38]. However, $\kappa_{\mathrm{peak}}$ does depend on the energy spectrum, e.g. via the slope $s$ in case of a power-law spectrum $\sim k^{-s}$. Flatter spectra implicate more energy on small scales, resulting in more contributions from high curvatures and consequently larger $\kappa_{\mathrm{peak}}$. This connection is illustrated in Figures 4, where $p_{\mathrm{MHD}}(\kappa)$ is modeled as a weighted sum of Gaussian components, similar to the description of the turbulent velocity increment distribution function as a Gaussian scale mixture [47, 48].
@@
4 Conclusion
We have presented a novel algorithm for synthetic magnetic turbulence based on a combination of the continuous cascade model, generalized to three-dimensional divergence-free vector fields, and the minimal multiscale Lagrangian map. The most important differences to previous works on the MMLM procedure are the explicit cascade structure of the underlying noise, primarily working with the vector potential, and interpolating from deformed to uniform grid once at the end of the procedure instead of interpolating at each scale. These changes considerably strengthen the emergent advective structures. We compare this algorithm with an incompressible resistive MHD turbulence simulation and the pure three-dimensional continuous cascade model, which is intermittent but lacks coherent structures. This comparison is done by means of visual inspection, the energy spectrum, conditional structure function scaling, the fieldline curvature distribution and running diffusion coefficients of charged test particles.